OpenStudy (anonymous):

solve (2b^2/-y^5)^-2

OpenStudy (anonymous):

i got y^10/4b^4 is this right

OpenStudy (zzr0ck3r):

solve?

OpenStudy (anonymous):

simplify sorry

OpenStudy (zzr0ck3r):

\[(\frac{2b^2}{-y^5})^{-2}=\frac{(2b^2)^{-2}}{(-y^5)^{-2}}=\frac{(-y^5)^2}{(2b^2)^2}=\frac{-y^{10}}{4b^4}\]

OpenStudy (zzr0ck3r):

yeah that shuold be (-y)^10 and that = y^10

OpenStudy (anonymous):

oh i forgot the negative thanksihaveon e more i'm not sure about

OpenStudy (zzr0ck3r):

\[(-y^5)^2=(-1*y^5)^2=(-1)^2(y^{10})=y^2\]

OpenStudy (zzr0ck3r):

err y ^10...

OpenStudy (zzr0ck3r):

you are correct:)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thanks

OpenStudy (zzr0ck3r):

np

OpenStudy (anonymous):

i have one more to show

OpenStudy (anonymous):

x-y^-1/x^-1-y ig ot xy-x-y^2/y is this correct?

OpenStudy (zzr0ck3r):

\[\frac{x-y^{-1}}{x^{-1}-y}?\]

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

\[\frac{x-\frac{1}{y}}{\frac{1}{x}-y}=\frac{\frac{xy-1}{y}}{\frac{1-xy}{x}}=\frac{xy-1}{y}*\frac{x}{-(xy-1)}=-\frac{x}{y }\]

OpenStudy (zzr0ck3r):

sorry I got a call...

OpenStudy (zzr0ck3r):

this is true for\[xy\ne1,x\ne0\]

OpenStudy (anonymous):

okay